OpenStudy (anonymous):

Find the probability of throwing a sum of 11 at least 8 times in 11 throws of a pair of fair dice.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (dumbcow):

binomial distribution \[P(x=k)=\left(\begin{matrix}n \\ k\end{matrix}\right)p^{k} (1-p)^{n-k}\] n=11 p = 1/18 (probability of rolling an 11) k = 8,9,10,11 answer is sum of probabilities for each "k" value